what re all the possible answers to the problem: 7 to the 8th power times 7 to the negative fourth power?
Answer by stanbon(75887)

what are all the possible answers to the problem: 7 to the 8th power times 7 to the negative fourth power?
------
7^8 *7^-4 = 7^8/7^4 = 7^(8-4) = 7^4
----------------------
Cheers,
Stan H.
------------